Anton McClure

Location: Akron, OH, US

                    E-Mail addresses: anton@next.gen.oh.us, asm@gnu.org
                     Web site: https://www.next.gen.oh.us/home/anton/

Free software / open-source seasoned university graduate seeks system administration career opportunities, with several years of experience, and a special interest in both cloud/datacenter-based and on-premise computing.

Skills

Programming languages: C, VB.Net, C++. Scripting languages: Perl, PHP, SH, CSH, KSH, BASH, JavaScript, QML, Python, PowerShell. Web servers: Apache (2.2 and 2.4), NGINX, Internet Information Services (IIS), lighttpd, LightSpeed Web Server, thttpd, Abyss Web Server, httpi. E-Mail servers: Sendmail, Exim 4, Courier, Postfix, Dovecot, Microsoft Exchange, Microsoft 365, Axigen, hMailServer. DNS servers: ISC BIND 9 (NAMED). Web development (coded): HTML/HTML5, XHTML/XHTML5, Server-Side Includes, CSS/CSS3, CGI, Custom CMS software. Web development (other): SEO, Google Analyrics, Microsoft Clarity, Oracle AddThis. Document typesetting: SGML, XML, LaTeX. Databases: MySQL/MariaDB, SQLite, PostgreSQL, MongoDB, Microsoft SQL, Microsoft Access. Microsoft Windows: System administration and support skills for Windows servers and clients. GNU/Linux & UNIX: System administration and support skills with GNU/Linux, FreeBSD, NetBSD, OpenBSD, and Sun/Oracle Solaris (including OpenSolaris & OpenIndiana). Platforms & services: Oracle PeopleSoft, Microsoft Azure, Microsoft Intune/Endpoint/Entra, Microsoft 365 Security, University of Akron’s UAnet Services (Auth). Software & tools: Microsoft Visual Studio, CVS, RCS, Git, Make, Microsoft 365, GNU GCC, Clang, OpenLDAP. Hardware: Repairing and replacing hardware, cleaning devices, maintaining RAID arrays, and re-imaging servers, clients, and other devices.

Work

Student Assistant, College of Business, University of Akron (August 2024-Present)

https://www.uakron.edu/cba/

• A student assistant at the University of Akron’s College of Business.

• Managed the college’s computer labs.

• Assisted in developing games for students at the Taylor Institute of Direct Marketing in their down time.

• Managed digital signage devices and software maintenance/updates throughout the College of Business building.

Lead Maintainer, GNU, FSF (February 2024-Present)

https://www.gnu.org/software/gnutrition/

• The Lead Maintainer of GNUtrition.

• Rewrote the program to bring GNUtrition to more people, remove the Python 2 and built-in GTK dependencies, and make the overall software more useful to more people.

• Wrote comprehensive documentation utilizing the GNU documentation standards.

Student Assistant, Information Technology Services, University of Akron (January 2023-May 2024)

https://www.uakron.edu/its/

• A student assistant at the University of Akron’s Information Technology Service Desk (ITSD). Started as an intern.

• Performed phone and remote technical support, computer repair, customer service, and other duties as assigned.

Projects

Next (2014-Present)

https://www.next.gen.oh.us/

• A group of computing services initially started in 2014 on a home server.

• Services provided includes a shell server, email server, authoritative DNS name servers and DNS resolvers, web server and web application hosting, among others.

• Main system runs the FreeBSD Operating System. Some other smaller machines run Debian, Ubuntu, and Rocky GNU/Linux.

Education

College of Engineering and Polymer Science, University of Akron (May 2023-Present)

https://www.uakron.edu/computer/cis/

• A Bachelor’s of Science in Computer Information Systems for the programming degree option.

• Currently attending for this degree.

College of Engineering and Polymer Science, University of Akron (Jan 2019-May 2023)

https://www.uakron.edu/computer/cis/

• An Associate of Applied Business degree in Computer Information Systems for the programming degree option.

• Graduated with degree in May of 2023.

Buchtel College of Arts and Sciences, University of Akron (Jan 2018-Jan 2019)

https://www.uakron.edu/computer/cs/

• A Bachelor’s of Science in Computer Science for the systems degree option.

• Transferred to associate degree for Computer Information Systems in December of 2018.

Links to latest version may be found at https://www.next.gen.oh.us/home/anton/professional/#vita.

Last updated on May 2, 2025.